### v1

@@ 3 Training
Our main goal was to train a highly performing, yet resource-friendly bilingual model while optimizing performances across both languages. To focus on the specific challenges of the bilingual paradigm, we often relied on previous work success to motivate our design and hyperparameter choices (Touvron et al., 2023b).
@@ 3.1 Model Architecture
We use the Llama architecture (Touvron et al., 2023a), a decoder-based transformer with a strong industrial focus. This model is trained with rotary position encodings and a context length of 2048. We construct 4 different model sizes by jointly scaling the number of attention heads, hidden size, and hidden layers. Table 2 summarizes the sizes of each model in the family.
@@ 3.2 Tokenizer
Most LLM tokenizers are fitted on English-centric corpora with an information-theoretic optimization objective, for example Byte-Pair encoding (Sennrich et al., 2016) or Unigram (Kudo, 2018), leading to good fertility values (low token per word ratio) on English text, but high fertility in other languages. These phenomena make processing in other languages slower and more costly. Furthermore, subword splits in non-English languages mechanically carry less semantical meaning, potentially being a factor in the degraded performance of models on non-English languages (Rust et al., 2021).
Tokenizer training. We fit our CroissantLLM tokenizer on a corpus of 100B high-quality tokens, with splits of English, French, and code data. We use Sentenpiece [^44] https://github.com/google/sentencepiece to train a Byte-Pair Encoding tokenizer with a vocabulary size of 32000 tokens, 100 special placeholder tokens, whitespace separation, and byte fallback, inspired by (Touvron et al., 2023a; Jiang et al., 2023). The data corpus used to fit the tokenizer is made available, [^45] https://huggingface.co/datasets/manu/tok-corpus-shuffled and notably contains large amounts of French data to skew the vocabulary construction process towards optimizing for French as well.
Improved fertility rates. The focus on English, French, and Code enables the CroissantLLM tokenizer to display smaller fertility rates on French texts than to the Mistral and Llama models with similar vocabulary sizes, all the while slightly outperforming them in English and Code (Figure 2). This is explained by the multilingual support of both Llama and Mistral tokenizers which need to allocate some vocabulary tokens to frequent character patterns from other languages. Roughly, the Llama tokenizer is 17 % less token efficient at encoding French internet data, and up to 40 % less efficient on clean encyclopedia French texts, implying the 303B unique French tokens in our data training set correspond to more than 360B tokens with the Llama tokenizer. This enables us to pack more data in fewer tokens, leading to improvements in training and inference efficiency.
@@ 3.3 Selecting an optimal language ratio
A crucial question when training a bilingual model is how to effectively weight data from the two languages during training to achieve a good trade-off between performance in both. While, intuitively, training on an equal mixture of English and French data may seem to be the obvious solution, differences in data quality available for each language coupled with transfer learning dynamics between both languages could imply that a balanced mix might be sub-optimal. However, training multiple models with different data mixes for comparison is prohibitively expensive.
@@
As such, we fit a joint scaling as described by Fernandes et al. (2023) for each language, by training 3 smaller model sizes on 3 different data mixes with varied ratios of English and French data (keeping the amount of Code data fixed). The corpus for these scaling law experiments is a subsampled variant of the larger corpus and is detailed in Appendix B.5. We define 3 data mixes by varying the language sampling ratio: (1) equal containing 40 % English data, 40 % French data and 20 % Code data; (2) frplus containing 20 % English data, 60 % French data and 20% Code data; and (3) enplus containing 60 % English data, 20 % French data and 20 % Code data. We then trained a 1.3B model on these subsets of the data for one of the data mixes to validate their predictive power.
@@ 3.5 Training framework
To maximize efficiency, we set the micro-batch size per device to 8 sequences of length 2048, and use 4 gradient accumulation steps, resulting in a total batch size of $8\times 4\times 30\times 8=7680$ samples, or $7680*2048=15728640$ tokens. We achieve a mean efficiency of 160 TFLOP/s, leading to a total compute estimate of $227.08e$ 22 FLOPS. Standard Cross-Entropy losses are used on a Causal Language Modeling objective.
@@ 3.6 Training losses
Training lasts 17 days for a total of 99648 GPU hours, and we chose not to manually intervene, letting the model recover on its own after occasional loss spikes. Curves suggest the model still has not reached a performance plateau after 3T tokens (Figure 5). Checkpoints are stored every 5k steps and released with the rest of the project artifacts.
@@ 3.7 Environmental impact
Interestingly, the model we trained is not “compute-optimal” according to Chinchilla laws (Hoffmann et al., 2022), meaning that less computing could have been used to train a larger model with the same performance. However, our model aims to be used for inference purposes at industrial scales. Our training paradigm is thus to absorb the downstream inference costs, by training a smaller model on a lot more tokens to obtain an inference-optimized model equivalent in performance to a bigger compute-optimal model (Sardana and Frankle, 2023). Each inference of the final model is thus vastly more energy-efficient than a Chinchilla optimal model of equivalent performance, and can even run on CPU or mobile devices. Relying on estimates of Kaplan et al. (2020b), at inference, CroissantLLM represents roughly 2.6 GFLOPS per token.
@@ 4 Evaluation Benchmarks
We hope to extend base model evaluation past English benchmarking alone and assess model capabilities in French, aiming for broad coverage across orthogonal capabilities to observe the effect of truly bilingual pre-training. Our evaluation efforts are rooted in transparency, and all results reported in the main technical report are reproducible through code that is open-sourced and public data. [^49] Evaluation is ran through the LM Evaluation harness (Gao et al., 2021) in a fork that will be merged upon acceptance https://github.com/ManuelFay/lm-evaluation-harness. Datasets used are grouped in the following collection: https://huggingface.co/collections/manu/frenchbench-evaluation-datasets-65bbcbf3987c4be8a5e9a2ef
@@ 4.2.1 FrenchBench Gen
FQuADQenQ is a question generation task in which passages and answers are given to the model in a few-shot manner, and we compute the ROUGE1 score (Lin, 2004) with the gold questions.
@@
MultiFQuAD  [^50] Previously unreleased dataset, evaluation set is released under CC-By-NC SA 4.0 license with this work is a FQuAD variant, with a publicly released evaluation set, in which answers can consist of multiple disjoint spans. We evaluate performance on the concatenation of these gold extractive spans using the ROUGE score.
@@ 4.2.2 FrenchBench Multiple Choice
French Hellaswag and Arc-E. These datasets correspond tp machine translations made by GPT3.5 of HellaSwag and Arc-C Easy to French. [^51] https://github.com/laiviet/lm-evaluation-harness/tree/main/datasets Manual verification of the translation quality indicates the translations to be far from perfect but sufficient for these datasets to act as a correct performance proxy.
OrangeSum. OrangeSum [^52] https://huggingface.co/datasets/orange_sum (Eddine et al., 2020) is a summarization dataset constructed from online News articles. Two standard French summarization tasks span from this dataset; OSum(T) in which the model is tasked with generating the title from the article body, and OSum(A) in which the model must generate the first paragraph of the article aimed to be an abstract of the article. Both tasks are measured through the ROUGE score.
@@ 4.3 Other Tasks
WMT14. Translation capabilities are evaluated through the test set of the 2014 WMT French-English and English-French tasks (Alves et al., 2023). We measure performance using BLEU score (sacreBLEU) (Papineni et al., 2002; Post, 2018), and COMET (Guerreiro et al., 2023b).
@@ 5 Benchmark results
For “monolingual” English models, we evaluate Pythia-1.4B (Biderman et al., 2023) trained on 300B tokens, OPT-1.3B (Zhang et al., 2022) trained on 180B tokens, and TinyLlama (1.1B) (Zhang et al., 2024). TinyLlama is a very strong English baseline, as it holds many similarities to CroissantLLM. It is a 1.1B model trained on 3 trillion tokens with the same English corpus as the Croissant base. Although it contains some amount of high-quality non-English data, it is only a minor share of the training corpus, the main data sources being English and code data. All models are trained way past Chinchilla optimality ($\sim$26B tokens for a 1.3B model).
For monolingual French models, we use GPT-fr (Simoulin and Crabbé, 2021), a 1B model trained on 16.3B tokens. We were not able to access the Pagnol model family (Launay et al., 2021) but show perplexity comparisons with reported results in Appendix D.2.
@@ 5.1 Base model
English. On English benchmarks (Table 3), CroissantLLM slightly trails behind TinyLlama, as is to be expected by the fact model capacity is split between two languages from our joint scaling laws experimentations. We however see training on such a large quantity of English tokens enables our model to edge out similarly sized monolingual models trained on fewer tokens (OPT, Pythia), and larger multiingual models (Bloom 3B) demonstrating the interest of pursuing training past Chinchilla optimality, especially when splitting model capacity across languages.
@@
CroissantCool 0.62 0.26 0.43 0.73 0.91 0.59
TinyLlama(1.1B) 0.65 0.26 0.45 0.73 0.94 0.60
Llama2(7B) 0.78 0.46 0.52 0.78 0.97 0.70
Mistral(7B) 0.83 0.85 0.55 0.81 0.98 0.80
@@
French. On French classification benchmarks, CroissantLLM logically largely outperforms models of similar sizes trained on mostly monolingual English or French data, and multilingual models (Table 4). Performance is on par with the Bloom(3B) model, which is about 3 times as large. An interesting phenomenon can be noticed, especially on generative benchmarks assessed in few-shot settings: “base” models trained with instruction-like data perform a lot better. This is noticeable with the Bloom(3B) model which outperforms the otherwise vastly superior Llama2(7B) model on several tasks, or through the performance gains of CroissantCool with respect to CroissantBase.
Task Hellaswag(fr) Arc-e(fr) fr-vocab fr-grammar Belebele(fr) Avg
@@
Table 4: FrenchBen MC (5-shot results)
Task FGenQ FGenAns MultiFQuAD OSum(A) OSum(T) Avg
mGPT(1.3B) 0.01 0.00 0.02 0.03 0.01 0.01
Pythia(1.4B) 0.13 0.00 0.00 0.00 0.00 0.03
GPT-fr(1B) 0.03 0.02 0.05 0.10 0.04 0.05
TinyLlama(1.1B) 0.03 0.14 0.12 0.10 0.08 0.09
CroissantLLM 0.13 0.23 0.21 0.11 0.11 0.16
Bloom(1.1B) 0.17 0.28 0.26 0.10 0.04 0.17
CroissantCool 0.17 0.31 0.28 0.20 0.13 0.22
Bloom(3B) 0.20 0.41 0.36 0.17 0.08 0.24
Llama2(7B) 0.22 0.39 0.26 0.16 0.20 0.25
Mistral(7B) 0.30 0.71 0.60 0.26 0.32 0.44
Table 5: FrenchBench Gen (5-shot results). Bloom models seem to have strong performance on QA tasks (Fquad), likely due to the inclusion of Question Answering datasets in its pretraining corpus (Laurençon et al., 2023).
Improvements throughout training. The model performance continues to improve on downstream tasks during the entirety of training. We report translation performance in Figure 6, and observe similar trends across all tasks. The benefits of training past Chinchilla optimality are clear, and although there are diminishing returns past a certain number of steps, training does not seem to saturate. In low training step settings, performance appears to emerge suddenly, reflecting emergent performance experiments in the literature most often obtained through model scaling (Wei et al., 2022b).
@@
Overall. The 1.1B CroissantLLM displays top-of-its-class performance across both languages and all benchmarks, even edging out larger models such as Bloom(3B) on most tasks. All models remain far off from the performance of the strong 7B Llama and Mistral models.
@@ 5.2.1 Chat Model
MT-Bench. We evaluate models on the MT-Bench benchmarks, both in English and French. Although a significant difference in performance can be noted between the Bloom model and Croissant in favor of the latter, performance differences with TinyLlama are not as important, neither in English nor in French. CroissantLLMChat performs strongly in open-ended writing categories (writing, roleplay, humanities) but struggles with reasoning and extractive tasks. Turn 2 performance (reformulation under constraints) is largely lower than Turn 1 performance as can be seen in Figures 11 and 12. Our CroissantLLMChat model also vastly outperforms the BloomZ 3B model trained by CMArkea on a large chat finetuning corpus (Delestre, 2023).
@@
French Trivia. One main question this work attempts to tackle is whether training on bilingual data goes beyond augmenting the language understanding and writing capabilities of a model in another language, but also equips the models with novel knowledge and different cultural biases. We evaluate French cultural knowledge on a Trivia task, consisting of questions about France-related topics, asked in English (Table 6). As heuristic-based metrics are insufficient to capture the diversity of possible answers (Faysse et al., 2023), we rely on GPT-4 to score the results. Similar results to the MT-Bench experiments can be noted: although CroissantLLM is the best performing, it is not so by a significant margin when compared to the best English model, TinyLlama, finetuned on the same bilingual SFT set. A potential bias might be that questions originate from American websites and asked in English, which might slightly favor TinyLlama. Both CroissantLLM and TinyLlama largely outperform Bloom pretrained on a smaller corpora but finetuned on the same data. This further underlines the importance of pretraining data quantity, and the large knowledge overlap between various sources and languages.
@@
Translation. We run translation evaluations on the Chat model and report results in Table 5.2.1. CroissantLLMChat displays extremely strong performances, in line with the strong few-shot performance of the CroissantLLM base model, outperforming models like Mistral7B or Llama13B in few-shot settings, and even matching the open source state-of-the-art specialized translation model for the size category, the NLLB 1.3B (Team et al., 2022), trained on vastly superior amounts of parallel data.
@@
Table 7: Performance in machine translation, according to COMET-22 and BLEU, across three different benchmarks: WMT14, TICO and FLORES. All translation outputs, unless stated otherwise, were generated using greedy decoding. We omit results with instructed models (–) on WMT14, since WMT14 was used during fine-tuning.
5.2.2 Summarization finetuning
To assess performance on specific downstream applications, we finetune base models on a dialog summarization dataset. [^55] Proprietary dataset belonging to Illuin Technology, corresponds to organic customer interactions with a counselor. Models are finetuned for three epochs on 6000 samples and results are computed through ROUGE and GPT-4 judgment metrics (Table 8).

### v4

@@ 3 Training
Our main goal was to train a highly performing, yet resource-friendly bilingual model while optimizing performances across both languages. To focus on the specific challenges of the bilingual paradigm, we rely on previous work to motivate many of our design and hyperparameter choices (Touvron et al., 2023b).
@@ 3.1 Model Architecture
We use the Llama architecture (Touvron et al., 2023a), a decoder-based transformer, trained with rotary position encodings (Su et al., 2023) and a context length of 2048. We construct 4 different model sizes by jointly scaling the number of attention heads, hidden size, and hidden layers. Table 2 summarizes the sizes of each model in the family.
@@ 3.2 Tokenizer
Most LLM tokenizers are fitted on English-centric corpora with an information-theoretic optimization objective, for example, Byte-Pair encoding (Sennrich et al., 2016) or Unigram (Kudo, 2018), leading to good fertility values (low token per word ratio) on English text, but high fertility in other languages. These phenomena make processing in other languages slower and more costly (Rust et al., 2021). Furthermore, subword splits in non-English languages mechanically carry less semantical meaning, potentially being a factor in the degraded performance of models on non-English languages (Rust et al., 2021).
Tokenizer training. We fit our CroissantLLM tokenizer on a corpus of 100B high-quality tokens, with splits of English, French, and code data. We use SentencePiece [^44] https://github.com/google/sentencepiece to train a Byte-Pair Encoding tokenizer with a vocabulary size of 32000 tokens, 100 special placeholder tokens, whitespace separation, and byte fallback, inspired by Touvron et al. (2023a); Jiang et al. (2023). The data corpus used to fit the tokenizer is made available, [^45] https://huggingface.co/datasets/manu/tok-corpus-shuffled and notably contains large amounts of French data to skew the vocabulary construction process towards optimizing for French as well.
Improved fertility rates. The focus on English, French, and Code enables the CroissantLLM tokenizer to display smaller fertility rates on French texts than the Mistral and Llama models with similar vocabulary sizes, all the while also displaying slightly smaller rates than both in English and Code (Figure 2). This is due to the multilingual support of both Llama and Mistral tokenizers which need to allocate some vocabulary tokens to frequent character patterns from other languages. Roughly, the Llama tokenizer is 17 % less token efficient at encoding French internet data, and up to 40 % less efficient on clean encyclopedia French texts, implying that the 303B unique French tokens in our data training set correspond to more than 360B tokens with the Llama tokenizer. This enables us to pack more data in fewer tokens, leading to improvements in training and inference efficiency.
@@ 3.3 Selecting an optimal language ratio
A crucial question when training a bilingual model is how to effectively weight data from the two languages to achieve a good trade-off between performance in both. While, intuitively, training on an equal mixture of English and French data may seem to be the obvious solution, differences in data quality available for each language coupled with transfer learning dynamics between both languages could imply that a balanced mix might be sub-optimal. However, training multiple models with different data mixes for comparison is prohibitively expensive.
@@
As such, we fit a joint scaling law as described by Fernandes et al. (2023) for each language, by training 3 smaller model sizes on 3 different data mixes with varied ratios of English and French data (keeping the amount of Code data fixed). The corpus for these scaling law experiments is a subsampled variant of the larger corpus and is detailed in Appendix B.5. We define 3 data mixes by varying the language sampling ratio: (1) equal containing 40 % English data, 40 % French data and 20 % Code data; (2) frplus containing 20 % English data, 60 % French data and 20% Code data; and (3) enplus containing 60 % English data, 20 % French data and 20 % Code data. We then trained a 1.3B model on these subsets of the data for one of the data mixes to validate their predictive power.
@@ 3.5 Training framework
To maximize efficiency, we set the micro-batch size per device to 8 sequences of length 2048, and use 4 gradient accumulation steps, resulting in a total batch size of $8\times 4\times 30\times 8=7680$ samples, or $7680*2048=15,728,640$ tokens. We achieve a mean efficiency of around 120 TFLOP484848160 TFLOP per second for our scaling law experiments with one GPU node only per second with activation checkpointing, leading to a total compute estimate of $224.30e$ 22 FLOPS. Standard Cross-Entropy losses are used on a Causal Language Modeling objective.
@@ 3.6 Training losses
Training lasts 17 days for a total of 99648 GPU hours, and we chose not to manually intervene, letting the model recover on its own after occasional loss spikes. We train with a max learning rate of $43e-$ 4, 1000 warmup steps, and a cosine learning rate with a minimum value of $51e-$ 5. Curves suggest the model still has not reached a performance plateau after 3T tokens (Figure 5). Checkpoints are stored every 5k steps and released with the rest of the project artifacts.
Refer to captionRefer to caption
@@ 3.7 Environmental impact
Interestingly, the model we trained is not “compute-optimal” according to Chinchilla laws (Hoffmann et al., 2022), meaning that less computing could have been used to train a larger model with the same performance. However, our model aims to be used for inference purposes at industrial scales. Our training paradigm is thus to absorb the downstream inference costs, by training a smaller model on a lot more tokens to obtain an inference-optimized model equivalent in performance to a bigger compute-optimal model (Sardana and Frankle, 2023). Each inference of the final model is thus vastly more energy-efficient than a Chinchilla optimal model of equivalent performance ($>$ 3B parameters), and can even run on CPU or mobile devices. Relying on estimates of Kaplan et al. (2020b), at inference, CroissantLLM represents roughly 2.6 GFLOPS per token.
@@ 4 Evaluation Benchmarks
We hope to extend base model evaluation past English benchmarking alone and assess model capabilities in French, aiming for broad coverage across orthogonal capabilities to observe the effect of truly bilingual pre-training. Our evaluation efforts are rooted in transparency, and all results reported in the main technical report are reproducible through code that is open-sourced and public data. [^50] Evaluation is ran through the LM Evaluation harness (Gao et al., 2021) at https://github.com/EleutherAI/lm-evaluation-harness. Datasets used are grouped in the following collection: https://huggingface.co/collections/manu/frenchbench-evaluation-datasets-65bbcbf3987c4be8a5e9a2ef
@@ 4.2.1 FrenchBench Gen
FQuADGenQ is a question generation task in which passages and answers are given to the model in a few-shot manner, and we compute the ROUGE1 score (Lin, 2004) with the gold questions.
@@
MultiFQuAD  [^51] Previously unreleased dataset, evaluation set is released under CC-By-NC SA 4.0 license with this work is a FQuAD variant, with a publicly released evaluation set, in which answers can consist of multiple disjoint spans. We evaluate performance on the concatenation of these gold extractive spans using the ROUGE1 score.
@@ 4.2.2 FrenchBench Multiple Choice
French Hellaswag and Arc-C. These datasets correspond to machine translations made by GPT3.5 of HellaSwag and Arc-C to French. [^53] https://github.com/laiviet/lm-evaluation-harness/tree/main/datasets Manual verification of the translation quality indicates the translations to be far from perfect but sufficient for these datasets to act as a correct performance proxy.
OrangeSum. OrangeSum [^54] https://huggingface.co/datasets/orange_sum (Eddine et al., 2020) is a summarization dataset constructed from online News articles. Two standard French summarization tasks span from this dataset; OSum(T) in which the model is tasked with generating the title from the article body, and OSum(A) in which the model must generate the first paragraph of the article aimed to be an abstract of the article. We select the abstract generation task, and measure performance with the ROUGE1 score.
@@ 4.3 Other Tasks
Translation. Translation capabilities are evaluated through the test set of the 2014 WMT French-English and English-French tasks (Alves et al., 2023). We measure performance using BLEU score (sacreBLEU, Papineni et al., 2002; Post, 2018), and COMET (Rei et al., 2022a). We also report FLORES (Team et al., 2022) and TICO (Anastasopoulos et al., 2020) scores.
@@ 5 Benchmark results
For “monolingual” English models, we evaluate Pythia-1.4B (Biderman et al., 2023) trained on 300B tokens, OPT-1.3B (Zhang et al., 2022) trained on 180B tokens, and TinyLlama(1.1B) (Zhang et al., 2024). TinyLlama is a very strong English baseline, as it holds many similarities to CroissantLLM. It is a 1.1B model trained on 3 trillion tokens with the same English corpus as the Croissant base. Although it contains some amount of high-quality non-English data, it is only a minor share of the training corpus, the main data sources being English and code data. As such, it trains on much more English tokens than CroissantLLM. All models are trained way past Chinchilla optimality ($\sim$26B tokens for a 1.3B model).
For monolingual French models, we use GPT-fr (Simoulin and Crabbé, 2021), a 1B model trained on 16.3B tokens, as well as the PagnolXL(1.5B) model (Launay et al., 2021), both in their author submitted HuggingFace implementations.
@@ 5.1 Base model
English. On English benchmarks (Table 3), CroissantLLM displays performances almost equivalent to those of TinyLlama, which has trained on much more English data. We see training on such a large quantity of English tokens enables our model to edge out similarly sized monolingual models trained on fewer tokens (OPT, Pythia), and larger multiingual models (Bloom 3B) demonstrating the interest of pursuing training past Chinchilla optimality, especially when splitting model capacity across languages.
@@
CroissantCool 0.62 0.26 0.43 0.73 0.92 0.59
TinyLlama(1.1B) 0.65 0.26 0.45 0.73 0.94 0.61
Llama2(7B) 0.79 0.46 0.56 0.79 0.97 0.72
Mistral(7B) 0.83 0.85 0.60 0.82 0.98 0.81
@@
French. On French classification benchmarks, CroissantLLM largely outperforms models of similar sizes trained on mostly monolingual English or French data, and multilingual models (Table 4). Performance is on par with the Bloom(3B) model, which is about 3 times as large. An interesting phenomenon can be noticed, especially on generative benchmarks assessed in few-shot settings: “base” models trained with instruction-like data perform a lot better. This is noticeable with the Bloom(3B) model which outperforms the otherwise vastly superior Llama2(7B) model on several tasks, or through the performance gains of CroissantCool with respect to CroissantBase.
Task Hellaswag(fr) Arc-c(fr) fr-vocab fr-grammar Belebele(fr) Avg
@@
Table 4: FrenchBench MC (5-shot results)
Task FGenQ FGenAns MultiFQuAD OSum(A) FTrivia Avg
Pagnol-XL(1.5B) 0.06 0.04 0.03 0.03 - 0.04
GPT-fr(1B) 0.04 0.02 0.05 0.11 - 0.06
mGPT(1.3B) 0.01 0.00 0.02 0.03 0.33 0.08
OPT(1.3B) 0.09 0.18 0.21 0.17 0.39 0.21
Bloom(1.1B) 0.17 0.28 0.26 0.10 0.31 0.23
Pythia(1.4B) 0.15 0.34 0.27 0.21 0.44 0.28
CroissantLLM 0.19 0.40 0.33 0.10 0.52 0.31
Bloom(3B) 0.21 0.47 0.37 0.18 0.47 0.34
TinyLlama(1.1B) 0.18 0.46 0.41 0.23 0.45 0.35
CroissantCool 0.20 0.45 0.36 0.27 0.53 0.36
Llama2(7B) 0.25 0.68 0.60 0.30 0.70 0.50
Mistral(7B) 0.33 0.78 0.64 0.31 0.74 0.56
Table 5: FrenchBench Gen (5-shot ROUGE1 results). Bloom models seem to have strong performance on QA tasks (Fquad), likely due to the inclusion of Question Answering datasets in its pretraining corpus (Laurençon et al., 2023). Pagnol-XL and GPT-fr are trained exclusively on French text and as such cannot be fairly evaluated on the French Trivia test.
Improvements throughout training. The model performance continues to improve on downstream tasks during the entirety of training. We report WMT14 translation performance in Figure 6, and observe similar trends across all tasks. The benefits of training past Chinchilla optimality are clear, and although there are diminishing returns past a certain number of steps, training does not seem to saturate. In low training step settings, performance appears to emerge suddenly, reflecting emergent performance experiments in the literature most often obtained through model scaling (Wei et al., 2022b).
@@
Overall. The 1.3B CroissantLLM displays top-of-its-class performance across both languages and all benchmarks, even edging out larger models such as Bloom(3B) on most tasks. All models remain far off from the performance of the strong 7B Llama and Mistral models.
@@ 5.2.1 Chat Model
MT-Bench. We evaluate models on the MT-Bench benchmarks, both in English and French. Although a large difference in performance can be noted between the Bloom model and Croissant in favor of the latter, performance differences with TinyLlama are not as significant, neither in English nor in French. CroissantLLMChat performs strongly in open-ended writing categories (writing, roleplay, humanities) but struggles with reasoning and extractive tasks. Turn 2 performance (reformulation under constraints) is largely lower than Turn 1 performance as can be seen in Figures 11 and 12. Our CroissantLLMChat model also vastly outperforms the BloomZ 3B model trained by CMArkea on a large chat finetuning corpus (Delestre, 2023).
@@
French Trivia. One main question this work attempts to tackle is whether training on bilingual data goes beyond augmenting the language understanding and writing capabilities of a model in another language, but also equips the models with novel knowledge and different cultural biases. We evaluate French cultural knowledge on a Trivia task, consisting of questions about France-related topics, asked in English (Table 6). As heuristic-based metrics are insufficient to capture the diversity of possible answers (Faysse et al., 2023), we rely on GPT-4 to score the results. Similar results to the MT-Bench experiments can be noted: although CroissantLLM is the best performing, no huge margins exist with respects to the best English model, TinyLlama, finetuned on the same bilingual SFT set. A potential bias might be that questions originate from American websites and are asked in English, which might slightly favor TinyLlama. Both CroissantLLM and TinyLlama largely outperform Bloom pretrained on a smaller corpora but finetuned on the same data. This further underlines the importance of pretraining data quantity, and the large knowledge overlap between various sources and languages.
@@
Translation. We run translation evaluations on the Chat models [^56] TinyLLaMAChat corresponds to the TinyLlama model finetuned with the same SFT set as CroissantLLMChat. and report results in Table 5.2.1. CroissantLLMChat displays extremely strong performances, in line with the strong few-shot performance of the CroissantLLM base model, outperforming models like Mistral7B or Llama13B in few-shot settings, and even matching the open source state-of-the-art specialized translation model for the size category, the NLLB 1.3B (Team et al., 2022), trained on vastly superior amounts of parallel data.
@@
Table 7: Performance in machine translation, according to COMET-22 and BLEU, across three different benchmarks: WMT14, TICO and FLORES. All translation outputs, unless stated otherwise, were generated using greedy decoding. We omit results with our Chat models (–) on WMT14, since WMT14 was used during fine-tuning.
5.2.2 Dialog Summarization finetuning
To assess performance on specific downstream applications, we finetune base models on a custom dialog summarization dataset. [^57] Proprietary dataset belonging to Illuin Technology, corresponds to organic customer interactions with a counselor. Models are finetuned for three epochs on 6000 samples and results are computed through ROUGE and GPT-4 judgment metrics (Table 8).